MONT 104N – Modeling the Environment Study Questions on “On Cannibals,” Michel de Montaigne August 24, 2012

1. Where did the word “barbarian” originally come from? What does Montaigne say about the way people tend to use the terms “barbaric” and “barbarian?” What examples does he give to demonstrate that the sort of “knee-jerk” reaction involved in declaring something unfamiliar to be “barbaric” can be mistaken?

2. What is the purpose of the long digression in Montaigne’s essay dealing with stories about the “lost continent” of Atlantis? Why does he include that whole discussion?

3. How does Montaigne contrast the “natural” and the “artificial?” Which is superior for him, and why? Do you agree with his point of view about that?

4. The main discussion in Montaigne’s essay concerns aspects of the culture of a group of natives of what is now Brazil who came into contact with Europeans during Montaigne’s lifetime. How does Montaigne describe their society? How does it compare with European society at the time Montaigne was writing? Would you say Montaigne expresses undiluted sympathy for and understanding of their customs? If not, what is his attitude?

5. What does Montaigne say about the ethical system of these Brazilian natives? What are the chief virtues they aim to exhibit? Why do they practice warfare?

6. For the Europeans who came in contact with these people, the most shocking aspect of their society was probably their practice of cannibalism. But exactly what function does cannibalism play in their culture? Why do they do this and how does Montaigne compare what they did to what Europeans did to prisoners of war? Is the practice of cannibalism justifiable in some circumstances?

7. Toward the end of the essay, Montaigne recounts the reactions of several members of this society who encounter the court of the French king Charles IX. What were there reactions to what they saw? Would you say they were being “barbaric?”

8. What does the last sentence of the essay mean? What attitude is Montaigne showing here?

9. Can we learn anything from Montaigne today? Why did we ask all students in the Holy Cross class of 2016 to read about him?